# Completeness conditions for spacetimes with low-regularity metrics

### v1

@@ Appendix A Examples of $C^{1}$-Spacetimes with geodesic uniqueness
In this appendix, we introduce examples of $C^{1}$-spacetimes where the uniqueness of causal geodesics holds and the exponential map as in Definition 4.8 can be defined.
@@
Let $g$ be as in (A.1). For any initial data
$\gamma(0)=(x^{0}_{0},x^{1}_{0}),\qquad\dot{\gamma}(0)\coloneqq(\tau_{0},\xi_{0})=(\dot{x}^{0}_{0},\dot{x}^{1}_{0})\in T_{(x^{0}_{0},x^{1}_{0})}\mathbf{R}^{2}$
with $\tau_{0}\neq 0$, there exist $\delta>0$ and a unique geodesic
@@
solving the geodesic equation with these initial conditions. In particular, every non-trivial causal geodesic satisfies $\tau_{0}\neq 0$, hence the above local uniqueness applies to all causal geodesics.
@@
Proposition A.2.
Under (A.1), $(\mathbf{R}^{2},g)$ is causally non-branching, and in particular maximally causally non-branching (MCNB) in the sense of Definition 4.5.
@@
Let $\gamma:[\alpha,\beta]\to\mathbf{R}^{2}$ be a non-trivial causal geodesic. As observed above, $\dot{x}^{0}(s)\neq 0$ for every interior point $s\in(\alpha,\beta)$. Fix such $s_{0}$. The initial data $(\gamma(s_{0}),\dot{\gamma}(s_{0}))$ then satisfy $\tau_{0}\coloneqq\dot{x}^{0}(s_{0})\neq 0$, so by Proposition A.1 the geodesic is locally unique. Hence any causal geodesic that agrees with $\gamma$ on $(s_{0}-\varepsilon,s_{0}]$ must coincide with it on $(s_{0},s_{0}+\delta)$, ruling out branching at interior points. Therefore $(\mathbf{R}^{2},g)$ is causally non-branching, and the same holds for maximal causal geodesics (MCNB). ∎
Theorem A.3.
If $\gamma\colon I\to\mathbf{R}^{2}$ and $\sigma\colon J\to\mathbf{R}^{2}$ are maximal causal geodesics for $g$ with
(A.13) $\gamma(0)=\sigma(0),\qquad\dot{\gamma}(0)=\dot{\sigma}(0)\neq 0,$
then $\gamma$ and $\sigma$ coincide on their maximal domains of definition.
Proof.
Proposition A.1 gives local agreement near $s=0$. By Proposition A.2, [KOSS22, Lemma 3.3] applies and this local agreement to coincidence on the entire maximal domains. ∎
Moreover, for the following class of $C^{1}$-spacetime which is causally non-branching and proper, the exponential map is $C^{2}$ on the future causal cone.
@@
Then for every $q\in M$, the exponential map $\exp_{q}$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$.
@@
with fixed $t_{*}\in\mathbf{R}$. Since $a\in C^{1}$ and $a\geq\alpha>0$ by (H1), we have $\sqrt{a}\in C^{1}$ and $\sqrt{a}>0$, hence $T\in C^{2}$ is strictly increasing. Therefore $T:\mathbf{R}\to I$ is a $C^{2}$-diffeomorphism onto an open interval $I\subset\mathbf{R}$, with inverse $t=t(\tau)\in C^{2}$. By (H2), $T(t)\to+\infty$ as $t\to+\infty$, so $I$ is unbounded above: there exists $A\in[-\infty,\infty)$ such that $I=(A,\infty)$.
Let $\Phi:\mathbf{R}^{2}\to I\times\mathbf{R}$ be the $C^{2}$-diffeomorphism $\Phi(t,x)=(\tau,x)=(T(t),x)$. In the $(\tau,x)$-coordinates, using $dt=d\tau/\sqrt{a(t(\tau))}$, we compute
@@
Fix $q\in M$, and set $Q\coloneqq\Phi(q)\in I\times\mathbf{R}$. For $v\in T_{q}M$, write $w\coloneqq d\Phi_{q}(v)\in T_{Q}(I\times\mathbf{R})\simeq\mathbf{R}^{2}$. If $v$ is future-directed nonzero causal, then in Minkowski coordinates $w=(w^{\tau},w^{x})$ satisfies $w^{\tau}>0$ and $(w^{\tau})^{2}\geq(w^{x})^{2}$. Along the $\eta$-geodesic with initial data $(Q,w)$, the solution is the straight line $s\mapsto Q+s\,w$. Since $I=(A,\infty)$ is unbounded above and $w^{\tau}>0$, the entire segment $\{\tau(Q)+sw^{\tau};s\in[0,1]\}$ lies in $I$, hence this geodesic exists on $[0,1]$. Therefore the Minkowski exponential is globally defined on the future causal cone and given by
@@
Here $v\mapsto d\Phi_{q}(v)$ is smooth, $w\mapsto Q+w$ is smooth, and $\Phi^{-1}$ is $C^{2}$. Consequently the composition $v\mapsto\exp_{q}^{g}(v)$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$. ∎

### v2

@@ Appendix A Examples of $C^{1}$-Spacetimes with geodesic uniqueness
In this appendix, we introduce examples of $C^{1}$-spacetimes where the uniqueness of causal geodesics holds and the causal exponential map as in Definition 4.8 can be defined.
@@
Let $g$ be as in (A.1). Fix an initial point $p\coloneqq(x^{0}_{0},x^{1}_{0})\in\mathbf{R}^{2}$ and an initial velocity $v\coloneqq(\tau_{0},\xi_{0})\in T_{p}\mathbf{R}^{2}$. If $\tau_{0}\neq 0$, there exist $\delta>0$ and a unique geodesic
@@
satisfying the initial conditions $\gamma(0)=p$ and $\dot{\gamma}(0)=v$. In particular, every non-trivial causal geodesic satisfies $\tau_{0}\neq 0$, hence the above local uniqueness applies to all causal geodesics.
@@
Theorem A.2.
Under the metric (A.1), $(M,g)$ is causally non-branching and non-intertwining.
@@
Let $\gamma$ be a non-trivial causal geodesic. Proposition A.1 establishes that for any point on $\gamma$, the solution to the geodesic equation is locally unique given the initial position and velocity.
This local uniqueness immediately implies the non-intertwining condition. Indeed, for any $p\in M$ and causal vector $v\in T_{p}M$, any two geodesics starting with $(p,v)$ must coincide on some time interval $0,\delta)$, which prevents them from intertwining.
Furthermore, since this local uniqueness holds at every point along any causal geodesic, branching cannot occur. Therefore, $(M,g)$ is causally non-branching. ∎
Remark A.3.
Theorem A.2 confirms that the spacetimes defined by (A.1) satisfy the assumptions required in Definition 4.8. Consequently, the causal exponential map is well-defined, and its continuity is guaranteed by Proposition 4.9.
As discussed above, any causally non-branching and non-intertwining $C^{1}$-spacetime admits a well-defined continuous causal exponential map, whereas for $C^{1,1}$-spacetimes, this map is a local bi-Lipschitz homeomorphism [KSS14, Theorem 2.1]. Moreover, the $C^{1}$-Lorentzian splitting theorem [BGMOS25, Theorem 2] establishes that a certain class of $C^{1}$-spacetimes is $C^{2}$-isometric to the product of $\mathbf{R}$ and a Riemannian manifold equipped with a $C^{1}$-metric. In the following theorem, we prove that for this specific class of spacetimes, the causal exponential map is $C^{2}$ on the future causal cone.
@@
Then for every $q\in M$, the causal exponential map $\exp_{q}$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$.
@@
with fixed $t_{*}\in\mathbf{R}$. Since $a\in C^{1}$ and $a\geq\alpha>0$ by (H1), we have $\sqrt{a}\in C^{1}$ and $\sqrt{a}>0$, hence $T\in C^{2}$ is strictly increasing. Therefore $T\colon\mathbf{R}\to I$ is a $C^{2}$-diffeomorphism onto an open interval $I\subset\mathbf{R}$, with inverse $t=t(\tau)\in C^{2}$. By (H2), $T(t)\to+\infty$ as $t\to+\infty$, so $I$ is unbounded above: there exists $A\in[-\infty,\infty)$ such that $I=(A,\infty)$.
Let $\Phi\colon\mathbf{R}^{2}\to I\times\mathbf{R}$ be the $C^{2}$-diffeomorphism $\Phi(t,x)=(\tau,x)=(T(t),x)$. In the $(\tau,x)$-coordinates, using $dt=d\tau/\sqrt{a(t(\tau))}$, we compute
@@
Fix $q\in M$, and set $Q\coloneqq\Phi(q)\in I\times\mathbf{R}$. For $v\in T_{q}M$, write $w\coloneqq d\Phi_{q}(v)\in T_{Q}(I\times\mathbf{R})\simeq\mathbf{R}^{2}$. If $v$ is future-directed nonzero causal, then in Minkowski coordinates $w=(w^{\tau},w^{x})$ satisfies $w^{\tau}>0$ and $(w^{\tau})^{2}\geq(w^{x})^{2}$. Along the $\eta$-geodesic with initial data $(Q,w)$, the solution is the straight line $s\mapsto Q+s\,w$. Since $I=(A,\infty)$ is unbounded above and $w^{\tau}>0$, the entire segment $\{\tau(Q)+sw^{\tau};s\in[0,1]\}$ lies in $I$, hence this geodesic exists on $[0,1]$. Therefore, the Minkowski exponential is globally defined on the future causal cone and given by
@@
Here $v\mapsto d\Phi_{q}(v)$ is smooth, $w\mapsto Q+w$ is smooth, and $\Phi^{-1}$ is $C^{2}$. Consequently, the composition $v\mapsto\exp_{q}^{g}(v)$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$. ∎